Nominal interest rate and extending credit

As a jewelry store manager, you want to offer credit, with interest on outstanding balances paid monthly. To carry receivables, you must borrow funds from your bank ata nominal 9%, monthly compounding. To offset your overhead, you want to charge your customers an EAR (or EFF%) that is 2% more than the bank is charging you. What APRrate should you charge your customers?
Answer #1
EAR = (1+i)^n - 1
SO Bank Borrowing has an EAR = (1+9%/12)^(1*12) - 1 =9.38% annual

So Owner should charges his customers an EAR = 9.38%+2% = 11.38% annual


answered by: Carlita